OpenStudy (anonymous):

Solve the equation. |-15| + 52 = 70y - |-33| - 69y

OpenStudy (anonymous):

15+52=70y-33-69y 67=y-33 y=100 Absolute value of a constant is always positive

OpenStudy (anonymous):

A. 73 B. -7 C. 7 D. -73